The following equations refer to the goods market of an economy in billions of euros:

C=480 +0,5Yd

I=110

T=70

G=250

1.Solve for goods market equilibrium 

2.Find equilibrium disposable income

3.find equilibrium consumption 

4.calculate the private savings,public savings and investment spending



1
Expert's answer
2020-08-18T13:07:07-0400

a)Y=C+I+G

Y=480+0.5(Y-70)+110+250

0.5Y=805

Y=1610

b)Yd=1610-70=1540

c)C=480+0.5(1540)=1250

d)S=Y-C

S=1610-1250=360